OpenStudy (anonymous):

would you set 16t^2-9 equal to 0 to factor as a trinomial?

OpenStudy (anonymous):

You will get (4t-3)(4t+3)

OpenStudy (anonymous):

do you mind drawing that out for me because I don't really understand how to get it?

OpenStudy (anonymous):

I did something wrong

OpenStudy (anonymous):

You sure the question says "equal to 0 to factor"

OpenStudy (anonymous):

it says factor

OpenStudy (mysesshou):

You did factor right, perhaps they meant to solve for t by factoring and setting equal to zero?

OpenStudy (anonymous):

It just says factor?

OpenStudy (anonymous):

Factor the expression

OpenStudy (anonymous):

Ok then I did do it right.

OpenStudy (mysesshou):

ok, then you don't need to set to zero to factor it, i suppose.

OpenStudy (mysesshou):

does the question have the zero and the trinomial part?

OpenStudy (anonymous):

just the trinomial

OpenStudy (mysesshou):

ok, then needhelpinmath first answer is right. you would only set it to zero if you were trying to solve for the t. This means (4t-3)=0 and (4t+3)=0 and solve for t. But if the question is really only asking for factor, then needhelpinmath's first answer is enough. :)

OpenStudy (anonymous):

thanks:]

OpenStudy (mysesshou):

\[16t^2-9\] Ones of this type follow the format \[(at-b)(at+b)=a^2t-b^2\] Here, the 16 gives \[16=a^2\] so the a=4. Then, the 9 gives \[9=b^2 \] so b=3. hope this helps some :)

OpenStudy (anonymous):

it does thank you very much

OpenStudy (mysesshou):

welcome ! :)

OpenStudy (mysesshou):

that was probably the process that needhelpinmath used to get to his/her factored solution.